# Ergodicity, KAM, FPUT

### v1

@@ 3 FPUT
In a sense the idea is that the Toda chains might play the role of an integrable Birkhoff normal form, of which fput-chains can be considered a perturbation. In AppendixD the basic results, [31], on the normal form of the periodic Toda-chains are summarized and heuristically applied to the fput-chains.
@@ Appendix A Shift equation
Remark the bound on the matrix elements $\sigma_{ij}=J_{0}^{-1}{\mbox{\boldmath$\partial$}}_{{\bf a}}^{2}{\overline{f}}_{0}({\bf a})$ in $C_{\lambda\rho_{0}}$ by $\sigma=((1-\lambda)\rho_{0})^{-1}J_{0,-}^{-1}\varepsilon_{0}$: hence the map $S{\bf a}={\bf a}+J_{0}^{-1}{\mbox{\boldmath$\partial$}}_{\bf a}{\overline{f}}({\bf a})$ has non singular Jacobian if $N\sigma<\frac{1}{3}$ (for instance)  141414 Let $|v|=\sum_{i}|v_{i}|$ for $v\in C^{N}$. Then the matrix $\delta_{ij}+\sigma_{ij}$ applied to $v\in C^{N}$ results in $w_{i}=v_{i}+\sum_{j}\sigma_{ij}v_{j}$: i.e. $|w|=\sum_{i}|w_{i}|\geq|v|(1-N\sigma)>0$ if, say, $N\sigma<\frac{1}{3}$.  and the map is locally invertible; furthermore any pair ${\bf a}^{\prime},{\bf a}^{\prime\prime}\in C((1-\lambda)\rho_{0})$ is mapped into a pair $S{\bf a}^{\prime},S{\bf a}^{\prime\prime}$ at distance $|S{\bf a}^{\prime}-S{\bf a}^{\prime\prime}|>(1-\sigma N)|{\bf a}^{\prime}-{\bf a}^{\prime\prime}|$, showing invertibility of $S$ in $C_{(1-\lambda)\rho_{0}}$ with image in $C((1+N\sigma)(1-\lambda)\rho_{0})\subset C(\rho_{0})$ if $\lambda$ is fixed $\lambda=\frac{2}{3}$ and $N\sigma<\frac{1}{3}$.
@@ Appendix B Renormalized H: dimensional estimates
Remark that in ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0})$ the Jacobian of the map ${\mbox{\boldmath$\alpha$}}^{\prime}={\mbox{\boldmath$\alpha$}}+{\partial}_{{{\bf A}}^{\prime}}\Phi_{0}$ is $(I+{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\alpha$}}{\mbox{\boldmath$\partial$}}_{{{\bf A}}^{\prime}}\Phi_{0})$ and, from the bound on $\Phi_{0}$, Eq.2.12 , $|{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\alpha$}}{\mbox{\boldmath$\partial$}}_{{{\bf A}}^{\prime}}\Phi_{0}|<6\widehat{g}\frac{\varepsilon_{0}}{C}\,\kappa_{0}^{-\widehat{c}}$.   [^15] Recall that $\Phi_{0}$ is bounded in ${\mathcal{C}}(\frac{2}{3}\rho_{0},\kappa_{0})$ and $|{\bf a}|<\frac{1}{6}\rho_{0}$; then use $\frac{2}{3}-\frac{1}{2}=\frac{1}{6}$. By Eq.2.12, $|{\partial}_{{{\bf A}}^{\prime}}\Phi_{0}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})|<\widehat{\gamma}\frac{\varepsilon_{0}}{C}\kappa_{0}^{-\widehat{c}}$ and the Jacobian determinant will be $\geq 1-\delta_{0}>0$ if $\delta_{0}\leq N!6\widehat{g}\frac{\varepsilon_{0}}{C}\,\kappa_{0}^{-\widehat{c}}$ is smaller than $1$: for definitness suppose
@@
Hence, for such $\delta_{0}$, the ${\mbox{\boldmath$\alpha$}}^{\prime}=S{\mbox{\boldmath$\alpha$}}={\mbox{\boldmath$\alpha$}}+{\mbox{\boldmath$\partial$}}_{{{\bf A}}^{\prime}}\Phi_{0}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$, maps (by Eq. 2.12) , ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0}-\delta_{0})$ into a set $\subset{\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0})$ covering the polydisk ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0}-2\delta_{0})$ and is also locally invertible (as its Jacobian is nonsingular) and globally one-to-one, hence holomorphic.  161616 Remark that, for the purpose of solving the implicit functions problems in Eq.2.4 , an argument can be used similar to the one for the solution of ${\bf a}=-J_{0}^{-1}{\mbox{\boldmath$\partial$}}_{\bf a}{\overline{f}}_{0}({\bf a})$ in Appendix A . Any pair ${\mbox{\boldmath$\alpha$}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime\prime}\in C(\frac{1}{2}\kappa_{0})$ is mapped, if $|v|=\sum_{i}|v_{i}|,v\in C^{N}$, into a pair $S{\mbox{\boldmath$\alpha$}}^{\prime},S{\mbox{\boldmath$\alpha$}}^{\prime\prime}$ at distance $|S{\mbox{\boldmath$\alpha$}}^{\prime\prime}-S{\mbox{\boldmath$\alpha$}}^{\prime}|>(1-N\|{\partial}_{{\mbox{\boldmath$\alpha$}}}{\partial}_{{{\bf A}}^{\prime}}\Phi_{0}\|_{\frac{1}{2}\rho_{0},\kappa_{0}})\,|{\mbox{\boldmath$\alpha$}}^{\prime}-{\mbox{\boldmath$\alpha$}}^{\prime\prime}|$; hence the map $S$ is one-to-one globally and analyticity of $S^{-1}$ will also follow from the positivity of the Jacobian. For a more general and systematic study of the inversion of the maps in Eq.2.4 see Appendicx G in [33]. $S^{-1}$ will be denoted ${\mbox{\boldmath$\alpha$}}={\mbox{\boldmath$\alpha$}}^{\prime}+\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$ and $\Delta$ inherits the bound Eq.2.13 $\|\Delta\|_{\frac{1}{2}\rho_{0},\kappa_{0}-2\delta_{0}}<{\overline{\gamma}}\frac{\varepsilon_{0}}{C}\,\kappa_{0}^{-{\overline{c}}}$ provided the conditions in Eq.(2.10) hold together with $\delta_{0}<\frac{1}{2}\kappa_{0}$, and the constants ${\overline{\gamma}},{\overline{c}}$ are chosen large enough. : its inverse in ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0}-2\delta_{0})$ will be written ${\mbox{\boldmath$\alpha$}}={\mbox{\boldmath$\alpha$}}^{\prime}+\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$ and fullfills the bound  2.13  for suitable ${\overline{\gamma}},{\overline{c}},\delta_{0}$.
@@
for $\gamma_{0},\widetilde{\gamma},\widetilde{c}$ suitable dimensionless constants.  [^17] Likewise the ${{\bf A}}={{\bf A}}^{\prime}+{\bf a}+{\partial}_{\mbox{\boldmath$\alpha$}}\Phi_{0}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$ would lead to ${{\bf A}}={{\bf A}}^{\prime}-{\bf a}+\Xi({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$ with $\Xi({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})=-{\partial}_{\mbox{\boldmath$\alpha$}}\Phi_{0}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}+\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}))$ under similar conditions; but $\Xi$ is not needed here.
Given a generic function $F({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$ define $F^{\prime}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}){\buildrel def\over{=}}F({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}+\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}))$. Then, by Eq.2.20, to estimate the derivatives ${\partial}_{{{\bf A}}^{\prime}}$ and ${\partial}_{{\mbox{\boldmath$\alpha$}}^{\prime}}$ of $F_{j}$ consider first the ${\partial}_{{{\bf A}}^{\prime}}$ and ${\partial}_{{\mbox{\boldmath$\alpha$}}}$. Then make use of, see also Eq. 2.102.13 :
${{\partial}_{{{\bf A}}^{\prime}}F^{\prime}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})=}$ ${{}{\partial}_{{{\bf A}}^{\prime}}F({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})+{\partial}_{\mbox{\boldmath$\alpha$}}F({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})\cdot{\partial}_{{{\bf A}}^{\prime}}\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})}$
${{\partial}_{{\mbox{\boldmath$\alpha$}}^{\prime}}F^{\prime}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})=}$ ${{}{\partial}_{{\mbox{\boldmath$\alpha$}}}F({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})\cdot(I+{\partial}_{{\mbox{\boldmath$\alpha$}}^{\prime}}\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}))}$
@@
${|{\partial}_{{{\bf A}}^{\prime}}\Delta|_{\frac{1}{4}\rho_{0},\kappa_{0}-2\delta_{0}}}$ ${{}+\frac{4}{\rho_{0}}|{\partial}_{{\mbox{\boldmath$\alpha$}}^{\prime}}\Delta|_{\frac{1}{4}\rho_{0},\kappa_{0}-2\delta_{0}}<{\overline{\gamma}}\frac{\varepsilon_{0}}{C}\kappa_{0}^{-{\overline{c}}}}$
Hence $f_{1}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})=f^{\prime}_{1}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})\equiv\sum_{j=1}^{3}F^{\prime}_{j}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$ and Eq.2.19:
${{}|{\partial}_{{{\bf A}}^{\prime}}f_{1}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})|=|{\partial}_{{{\bf A}}^{\prime}}f^{\prime}_{0}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})+{\partial}_{{\mbox{\boldmath$\alpha$}}}f^{\prime}_{0}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}){\partial}_{{{\bf A}}^{\prime}}\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})|}$
@@
${{}|\frac{1}{\widetilde{\rho}}{\partial}_{{\mbox{\boldmath$\alpha$}}^{\prime}}f^{\prime}_{1}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})|=|\frac{1}{\widetilde{\rho}}{\partial}_{{\mbox{\boldmath$\alpha$}}}f^{\prime}_{0}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})(I+{\partial}_{{\mbox{\boldmath$\alpha$}}^{\prime}}\Delta({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime}))|}$
${{}\leq|\frac{1}{\widetilde{\rho}}{\partial}_{{\mbox{\boldmath$\alpha$}}}f^{\prime}_{0}({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})|(1+N{\overline{\gamma}}\frac{\varepsilon_{0}}{C}\kappa_{0}^{-{\overline{c}}})}$

### v2

@@ 3 FPUT
In a sense the idea is that the Toda chains might play the role of an integrable Birkhoff normal form, of which fput-chains can be considered a perturbation. In the next section the basic results, [31], on the normal form of the periodic Toda-chains are summarized and heuristically applied to the fput-chains.
4 Toda and FPUT: heuristics
$H({\bf p},{\bf q})=\sum_{k=1}^{N}\Big(\frac{1}{2}p_{k}^{2}+\alpha^{2}e^{q_{k}-q_{k+1}}\Big)=\sum_{k=1}^{N}(\frac{1}{2}p_{k}^{2}+\alpha^{2}V_{T}(q_{k}-q_{k+1}))$ (4.1)
The periodic Toda lattice model is studied in [31] where $2N$ global canonical action-angle variables $({\bf a},{\mbox{$\varphi$}})\in(R_{+}^{N},[0,2\pi]^{N})$ are constructed and evolve at time $t$ into $({\bf a},{\mbox{$\varphi$}}+{\mbox{$\omega$}}({\bf a})t)$, following a Hamiltonian $H_{\alpha}({\bf a})$. Near ${\bf a}=0$ it is $\omega_{i}({\bf a})=2\alpha\omega_{0,i}+\frac{1}{4N}a_{i}+O(|{\bf a}|^{2})$, if ${\mbox{$\omega$}}_{0}$ is the spectrum of the linear oscillations $\omega_{0,i}=2\sin\frac{\pi i}{N}$.
The $H_{\alpha}({\bf a})$, [31, th.1.4], is globally defined for ${\bf a}\in R^{N}_{+}$ and the Jacobian $J_{\alpha}({\bf a})=\det\frac{{\partial}{\mbox{$\omega$}}({\bf a})}{{\partial}{\bf a}}$, 3, is analytic in $\alpha>0$; for $\alpha>0$ it is strictly convex in ${\bf a}\in R_{+}^{N}$ and $J_{\alpha}({\bf a})_{ij}=\frac{1}{4N}\delta_{ij}+O(|{\bf a}|)\neq 0$ and the spectrum is $\omega_{i}=2\alpha\sin(\frac{i\pi}{N})+\frac{1}{4N}a_{i}+O(|{\bf a}|^{2})$.
However application to the fput-chains requires more analysis mainly because the use of the change of variables leads to the Hamiltonian in Eq.4.1 with $\alpha$ small and compare it with Eq.3.6 $\varepsilon$ small: aside from the time rescaling, this means to compare motions with $\alpha$ in Eq.4.1 to motions with $\varepsilon$ in Eq.3.6 setting $\alpha=\varepsilon^{-2}$; so when $\alpha$ is very large.
The analysis in [31] is global and the coordinates $({\bf a},{\mbox{$\varphi$}})$ exist for all $\alpha>0$ and transform analytically the Hamiltonian Eq.4.1 into a function $\widetilde{H}_{\alpha}({\bf a})$ with convex Jacobian $J_{ij}=\frac{{\partial}^{2}_{a_{i}}H({\bf a})}{{\partial}^{2}a_{j}}$, see th.(1.1),th.(1.2) in [31]. Hence in the region of couplings relevant for the fput-chains, i.e. $\varepsilon$ close to 0, $\alpha=\varepsilon^{-2}$ large, the Toda lattice Hamiltonian admits action-angles coordinates.
Proceeding heuristically: the Toda lattice can be considered a perturbation of order $\varepsilon^{3}$ of the fput-lattice $H_{3,4}$, see Eq.3.6:
$H_{3}({\bf p},{\bf q})=\sum_{j=0}^{N-1}\Big(\frac{1}{2}(p_{k}^{2}+(q_{k}-q_{k+1})^{2})+\frac{\varepsilon}{3!}(q_{k}-q_{k+1})^{3}+\frac{\varepsilon^{2}}{4!}(q_{k}-q_{k+1})^{4}\Big)$ (4.2)
Hence the periodic elastic chain perturbed by $h_{3,4}=\frac{\varepsilon}{3!}\sum_{j=0}^{N-1}(q_{k}-q_{k+1})^{3}+\frac{\varepsilon^{2}}{4!}\sum_{j=0}^{N-1}(q_{k}-q_{k+1})^{4}$ and the periodic Toda lattice truncated to order $\varepsilon^{2}$ are the same: thus their first two orders Birkhoff’s normal forms, to remove the resonant terms present in $H_{3,4}$, coincide up to the next order above $\varepsilon^{2}$ (the first steps in the construction of Birkhoff normal form only involves the common part $h_{3,4}$). The integrability of the Toda lattice should then imply that the normal form for the above truncation is integrable implying that the normal form of the above fput-chain is integrable to order $\varepsilon^{2}$.
Therefore the argument of [26] can be applied (heuristically here, and avoiding any explict calculation of the Birkhoff normal form) to reach the same conclusion and the periodic fput-chain model in Eq.4.2 shows quasi periodic motions at small $\varepsilon$, filling large regions of phase space in any region $\Omega$ where $e<H<E$, up to a set of volume approching 0 as $\varepsilon\to 0$.
The chain considered in Eq.4.2, $H_{3,4}$, has a Birkhoff’s normal form which to order $\varepsilon^{2}$ is the sum of the (resonant) contributions to the $\alpha$-chain (to second order) and those to the $\beta$ chain (to first order): making use of the arbitrariness of $\varepsilon$ it follows that both $\alpha$ and $\beta$ chains have an integrable normal form, to order $2$ included, and the KAM theory can be applied to both, if the Jacobian $J$ 3 is non singular (as it is in the Toda lattice and in the $\alpha$-model, [31, 26]).
@@ Appendix A Shift equation
Remark the bound on the matrix elements $\sigma_{ij}=J_{0}^{-1}{\mbox{$\partial$}}_{{\bf a}}^{2}{\overline{f}}_{0}({\bf a})$ in $C_{\lambda\rho_{0}}$ by $\sigma=((1-\lambda)\rho_{0})^{-1}J_{0,-}^{-1}\varepsilon_{0}$: hence the map $S{\bf a}={\bf a}+J_{0}^{-1}{\mbox{$\partial$}}_{\bf a}{\overline{f}}({\bf a})$ has non singular Jacobian if $N\sigma<\frac{1}{3}$ (for instance)   [^14] Let $|v|=\sum_{i}|v_{i}|$ for $v\in C^{N}$. Then the matrix $\delta_{ij}+\sigma_{ij}$ applied to $v\in C^{N}$ results in $w_{i}=v_{i}+\sum_{j}\sigma_{ij}v_{j}$: i.e. $|w|=\sum_{i}|w_{i}|\geq|v|(1-N\sigma)>0$ if, say, $N\sigma<\frac{1}{3}$.  and the map is locally invertible; furthermore any pair ${\bf a}^{\prime},{\bf a}^{\prime\prime}\in C((1-\lambda)\rho_{0})$ is mapped into a pair $S{\bf a}^{\prime},S{\bf a}^{\prime\prime}$ at distance $|S{\bf a}^{\prime}-S{\bf a}^{\prime\prime}|>(1-\sigma N)|{\bf a}^{\prime}-{\bf a}^{\prime\prime}|$, showing invertibility of $S$ in $C_{(1-\lambda)\rho_{0}}$ with image in $C((1+N\sigma)(1-\lambda)\rho_{0})\subset C(\rho_{0})$ if $\lambda$ is fixed $\lambda=\frac{2}{3}$ and $N\sigma<\frac{1}{3}$.
@@ Appendix B Renormalized H: dimensional estimates
Remark that in ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0})$ the Jacobian of the map ${\mbox{$\alpha$}}^{\prime}={\mbox{$\alpha$}}+{\partial}_{{{\bf A}}^{\prime}}\Phi_{0}$ is $(I+{\mbox{$\partial$}}_{\mbox{$\alpha$}}{\mbox{$\partial$}}_{{{\bf A}}^{\prime}}\Phi_{0})$ and, from the bound on $\Phi_{0}$, Eq.2.12 , $|{\mbox{$\partial$}}_{\mbox{$\alpha$}}{\mbox{$\partial$}}_{{{\bf A}}^{\prime}}\Phi_{0}|<6\widehat{g}\frac{\varepsilon_{0}}{C}\,\kappa_{0}^{-\widehat{c}}$.   [^15] Recall that $\Phi_{0}$ is bounded in ${\mathcal{C}}(\frac{2}{3}\rho_{0},\kappa_{0})$ and $|{\bf a}|<\frac{1}{6}\rho_{0}$; then use $\frac{2}{3}-\frac{1}{2}=\frac{1}{6}$. By Eq.2.12, $|{\partial}_{{{\bf A}}^{\prime}}\Phi_{0}({{\bf A}}^{\prime},{\mbox{$\alpha$}})|<\widehat{\gamma}\frac{\varepsilon_{0}}{C}\kappa_{0}^{-\widehat{c}}$ and the Jacobian determinant will be $\geq 1-\delta_{0}>0$ if $\delta_{0}\leq N!6\widehat{g}\frac{\varepsilon_{0}}{C}\,\kappa_{0}^{-\widehat{c}}$ is smaller than $1$: for definitness suppose
@@
Hence, for such $\delta_{0}$, the ${\mbox{$\alpha$}}^{\prime}=S{\mbox{$\alpha$}}={\mbox{$\alpha$}}+{\mbox{$\partial$}}_{{{\bf A}}^{\prime}}\Phi_{0}({{\bf A}}^{\prime},{\mbox{$\alpha$}})$, maps (by Eq. 2.12) , ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0}-\delta_{0})$ into a set $\subset{\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0})$ covering the polydisk ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0}-2\delta_{0})$ and is also locally invertible (as its Jacobian is nonsingular) and globally one-to-one, hence holomorphic.   [^16] Remark that, for the purpose of solving the implicit functions problems in Eq.2.4 , an argument can be used similar to the one for the solution of ${\bf a}=-J_{0}^{-1}{\mbox{$\partial$}}_{\bf a}{\overline{f}}_{0}({\bf a})$ in Appendix A . Any pair ${\mbox{$\alpha$}}^{\prime},{\mbox{$\alpha$}}^{\prime\prime}\in C(\frac{1}{2}\kappa_{0})$ is mapped, if $|v|=\sum_{i}|v_{i}|,v\in C^{N}$, into a pair $S{\mbox{$\alpha$}}^{\prime},S{\mbox{$\alpha$}}^{\prime\prime}$ at distance $|S{\mbox{$\alpha$}}^{\prime\prime}-S{\mbox{$\alpha$}}^{\prime}|>(1-N\|{\partial}_{{\mbox{$\alpha$}}}{\partial}_{{{\bf A}}^{\prime}}\Phi_{0}\|_{\frac{1}{2}\rho_{0},\kappa_{0}})\,|{\mbox{$\alpha$}}^{\prime}-{\mbox{$\alpha$}}^{\prime\prime}|$; hence the map $S$ is one-to-one globally and analyticity of $S^{-1}$ will also follow from the positivity of the Jacobian. For a more general and systematic study of the inversion of the maps in Eq.2.4 see Appendicx G in [33]. $S^{-1}$ will be denoted ${\mbox{$\alpha$}}={\mbox{$\alpha$}}^{\prime}+\Delta({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})$ and $\Delta$ inherits the bound Eq.2.13 $\|\Delta\|_{\frac{1}{2}\rho_{0},\kappa_{0}-2\delta_{0}}<{\overline{\gamma}}\frac{\varepsilon_{0}}{C}\,\kappa_{0}^{-{\overline{c}}}$ provided the conditions in Eq.(2.10) hold together with $\delta_{0}<\frac{1}{2}\kappa_{0}$, and the constants ${\overline{\gamma}},{\overline{c}}$ are chosen large enough. : its inverse in ${\mathcal{C}}(\frac{1}{2}\rho_{0},\kappa_{0}-2\delta_{0})$ will be written ${\mbox{$\alpha$}}={\mbox{$\alpha$}}^{\prime}+\Delta({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})$ and fullfills the bound  2.13  for suitable ${\overline{\gamma}},{\overline{c}},\delta_{0}$.
@@
for $\gamma_{0},\widetilde{\gamma},\widetilde{c}$ suitable dimensionless constants.  [^17] Likewise the ${{\bf A}}={{\bf A}}^{\prime}+{\bf a}+{\partial}_{\mbox{$\alpha$}}\Phi_{0}({{\bf A}}^{\prime},{\mbox{$\alpha$}})$ would lead to ${{\bf A}}={{\bf A}}^{\prime}-{\bf a}+\Xi({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})$ with $\Xi({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})=-{\partial}_{\mbox{$\alpha$}}\Phi_{0}({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime}+\Delta({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime}))$ under similar conditions; but $\Xi$ is not needed here.
Given a generic function $F({{\bf A}}^{\prime},{\mbox{$\alpha$}})$ define $F^{\prime}({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime}){\mathrel{\mathop{\kern 0.0pt=}\limits^{def}}}F({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime}+\Delta({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime}))$. Then, by Eq.2.20, to estimate the derivatives ${\partial}_{{{\bf A}}^{\prime}}$ and ${\partial}_{{\mbox{$\alpha$}}^{\prime}}$ of $F_{j}$ consider first the ${\partial}_{{{\bf A}}^{\prime}}$ and ${\partial}_{{\mbox{$\alpha$}}}$. Then make use of, see also Eq. 2.102.13 :
${{\partial}_{{{\bf A}}^{\prime}}F^{\prime}({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})=}$ ${{}{\partial}_{{{\bf A}}^{\prime}}F({{\bf A}}^{\prime},{\mbox{$\alpha$}})+{\partial}_{\mbox{$\alpha$}}F({{\bf A}}^{\prime},{\mbox{$\alpha$}})\cdot{\partial}_{{{\bf A}}^{\prime}}\Delta({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})}$
${{\partial}_{{\mbox{$\alpha$}}^{\prime}}F^{\prime}({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})=}$ ${{}{\partial}_{{\mbox{$\alpha$}}}F({{\bf A}}^{\prime},{\mbox{$\alpha$}})\cdot(I+{\partial}_{{\mbox{$\alpha$}}^{\prime}}\Delta({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime}))}$
@@
${|{\partial}_{{{\bf A}}^{\prime}}\Delta|_{\frac{1}{4}\rho_{0},\kappa_{0}-2\delta_{0}}}$ ${{}+\frac{4}{\rho_{0}}|{\partial}_{{\mbox{$\alpha$}}^{\prime}}\Delta|_{\frac{1}{4}\rho_{0},\kappa_{0}-2\delta_{0}}<{\overline{\gamma}}\frac{\varepsilon_{0}}{C}\kappa_{0}^{-{\overline{c}}}}$
Hence $f_{1}({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})=f^{\prime}_{1}({{\bf A}}^{\prime},{\mbox{$\alpha$}})\equiv\sum_{j=1}^{3}F^{\prime}_{j}({{\bf A}}^{\prime},{\mbox{$\alpha$}})$ and Eq.2.19:
${{}|{\partial}_{{{\bf A}}^{\prime}}f_{1}({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})|=|{\partial}_{{{\bf A}}^{\prime}}f^{\prime}_{0}({{\bf A}}^{\prime},{\mbox{$\alpha$}})+{\partial}_{{\mbox{$\alpha$}}}f^{\prime}_{0}({{\bf A}}^{\prime},{\mbox{$\alpha$}}){\partial}_{{{\bf A}}^{\prime}}\Delta({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})|}$
@@
${{}|\frac{1}{\widetilde{\rho}}{\partial}_{{\mbox{$\alpha$}}^{\prime}}f^{\prime}_{1}({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})|=|\frac{1}{\widetilde{\rho}}{\partial}_{{\mbox{$\alpha$}}}f^{\prime}_{0}({{\bf A}}^{\prime},{\mbox{$\alpha$}})(I+{\partial}_{{\mbox{$\alpha$}}^{\prime}}\Delta({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime}))|}$
${{}\leq|\frac{1}{\widetilde{\rho}}{\partial}_{{\mbox{$\alpha$}}}f^{\prime}_{0}({{\bf A}}^{\prime},{\mbox{$\alpha$}})|(1+N{\overline{\gamma}}\frac{\varepsilon_{0}}{C}\kappa_{0}^{-{\overline{c}}})}$